Calculate the partial pressure of all the components in a mixture that contains 1 mole of CO2, 2.5 moles of CO, and 2.6 moles of CH4 in a constant volume vessel at 130 mm Hg and 300C


1
Expert's answer
2022-02-03T01:06:09-0500

Mole ratios= CO2: CO: CH4

= 1: 2.5: 2.6

Partial pressure of CO2=( 1/6.1)× 130= 21.3mmHg

Partial pressure of CO=( 2.5/6.1)× 130= 53.3mmhg

Partial pressure of CH4= (2.6/6.1)× 130= 55.4mmHg